OpenStudy (anonymous):

evaluate the function at the given values of x, if possible f(x)=2x^2-1/x2 (a)x=5 (b)x=-3

OpenStudy (anonymous):

kianap just replc the value instead of x ... I'm sure u will get the answer : )

OpenStudy (anonymous):

if u still have problem then pls tell me .. and is the denominator x^2 or just x2

OpenStudy (anonymous):

this problems is more difficult than the last one.

OpenStudy (anonymous):

ok , f(x) = 2x^2 - 1 -------- right? x^2

OpenStudy (anonymous):

is denominator \[2x\]\[x ^{2}\] ?

OpenStudy (anonymous):

the second one, sorry

OpenStudy (anonymous):

ok : )

jhonyy9 (jhonyy9):

for x=5 f(5)=50-1/25 for x=3 f(3)=18-1/9

OpenStudy (anonymous):

Now , f(x) = 2x^2 - 1 -------- x^2 our first value is 5 5^2 =25 so f(5) = 2*25 -1 /25 f(5) = 49/25

OpenStudy (anonymous):

for f(-3) = 2 *-3*-3 - 1 -------------- -3*-3 f(-3)= 17/9

OpenStudy (anonymous):

thank u... i am writting this example down but, i still really do not understand the square root... I will get it.